Fermi National Laboratory


Accelerator Update

Monday April 19
The day shift began with Tevatron (TeV) experts investigating separator problems, with the Antiproton Source (Pbar) stacking, with Operations monitoring a Recycler stash, and with MiniBooNE and the Meson experiment MCenter taking beam.

At 10 AM, Operations turned off beam to MiniBooNE while a Linac expert conducted beam diagnostics. MiniBooNE was turned back on at 10:23 AM.

A Booster expert began working on the Booster chopper at 10:24 AM. He finished his work an hour later.

Operations reported at 1:56 PM that the Main Injector RF pond water strainer at MI60 was clogged and technicians began cleaning it. This halted MI and Switchyard (SY) beam until around 3:49 PM.

Operations put the TeV through a wet squeeze at 3:52 PM and TeV experts continued with studies.

Around 4 PM, the TeV suffered a quench at sector A1. Operators believed that the quench had something to do with aborting beam on the central orbit.

Operations put the TeV through a dry squeeze at 6:37 PM.

At 8:19 PM, Operations put the TeV into shot setup.

At 9:49 PM, while loading the final protons into the TeV, the protons aborted due to CAMAC crate power supply failure at the F1 service building. Operators had the crate and power supply replaced by 10:40 PM.

Tuesday April 20
The midnight shift began with the TeV in shot setup, with MiniBooNE taking beam, and with MCenter in access.

At 12:51 AM, Operations established store 3421 with an initial luminosity of 37.38E30.

Operations was going to conduct a Recycler stash at 3 AM, but a RF problem put the stash on hold.

Pbar resumed stacking at 3:14 AM.

The Recycler stash began at 3:51 AM and ended with the fourth and last shot of antiprotons at 4:42 AM. Pbar resumed stacking.

A Booster RF station (BRF16) tripped off at 5:19 AM and then tripped off again 50 minutes later. Then at 6:11 AM, it tripped off for the first of seven times in eight minutes. Operators turned the station off at 6:20 AM.

Operations established beam to MCenter at 6:54 AM.

At 12:58 PM, a TeV RF station (TRF7) tripped off.

Pbar and MI experts began studies at 2:28 PM.

At 3:10 PM, nearly all of the MI RF stations tripped off. A MI expert got the stations back on. He found and repaired a bad cable.

MI and Pbar studies ended at 5:06 PM.

A MI RF station tripped off due to a fan problem at 6:49 PM. Operators bypassed the station.

Operations reported a power glitch at 7:21 PM that tripped off the Booster West Anode power supply and the Tevatron Electron Lens (TEL). The Collider Detector at Fermilab (CDF) and D-zero (D0) experiments went to standby until the TEL could be turned back on. Several Linac Klystron RF station tripped off also.

Pbar resumed stacking at 7:25 PM and a minute later the TEL came back on.

At 7:36 PM, a tornado warning caused all the operators to leave the Main Control Room (MCR) and go to the basement. They returned to the MCR twenty minutes later. CDF and D0 resumed normal operations.

Operations resumed sending beam to MCenter at 9 PM.

Wednesday April 21
The midnight shift began with Operations monitoring store 3421 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:33 AM, Operations terminated store 3421.

Operations reported a horizontal separator spark at the TeV sector B17 separator.

Operations put the TeV into shot setup at 2:40 AM.

At 5:12 AM, Operations established store 3423 with an initial luminosity of 47E30. Operations reestablished beam to MCenter.

Recycler began stashing antiprotons at 7 AM and ended at 8 AM.

Plans for Today
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/21/2004   email Fermilab